OpenStudy (anonymous):

Find an equation of the line satisfying the conditions given. Express your answer in standard form. Parallel to 3x - y = -5 and passing through ( -1, 0 )

OpenStudy (bahrom7893):

If it is parallel to the line 3x - y = -5, then it must have the same slope as that line: Let me rewrite your equation, so that it is easier for me to see the slope: 3x - y = -5, move 3x to the right hand side: - y = - 5 - 3x, multiply thru by -1: y = 5 + 3x or y = 3x +5. This is in the form of y = Mx + b, where M = slope = 3

OpenStudy (bahrom7893):

Okay so now the formula of the line is: \[y - y _{1} = M ( x - x _{1})\] In your case: \[x_{1} = -1\] and \[y_{1} = 0\] ( -1 ; 0 )

OpenStudy (bahrom7893):

so you know that M = 3; x1 = -1 and y1 = 0. Plug these into the equation of a line above: y - ( 0 ) = 3 * ( x - (-1) ) Simplify: y = 3x + 3 <=Your answer

OpenStudy (bahrom7893):

any questions?